OpenStudy (anonymous):

find the slope and the y-intercept of the equation y-3=5

OpenStudy (anonymous):

First solve for y. \[y - 3 = 5 \implies y = 8\]Now, this is in the form \(y = mx + b\). b is the y-intercept. m is the slope. Notice how the m disappeared. What times anything is 0?

OpenStudy (anonymous):

ok i thought the slope was 3 and the y intercepts were (0,5)??

OpenStudy (anonymous):

If you look farther into it: \[y = 0x + 8\]\[y = mx + b\]Do you see how the slope is 0 and how the y-intercept is (0, 8)?

OpenStudy (anonymous):

i'm a little confused, but i'm re- reading the lesson maybe i missed something

OpenStudy (anonymous):

Ok. When the x-term disappears and it's in the form y = ?, then it's a horizontal line with a slope of 0. When the y-term disappears and it's in the form x = ?, then it's a vertical line with a slope of undefined.

OpenStudy (anonymous):

ok I see now, thank you very much

OpenStudy (anonymous):

np :)